# FLEX: A Large-Scale Multi-Modal Multi-Action Dataset for Fitness Action Quality Assessment

### v1

@@ 3.2 Data Collection
Fitness actions involve people in convoluted positions and postures (not captured in typical pose estimation datasets). Additionally, in weight training exercises, essential body parts and joints might be occluded if the scene is captured from any single camera angle. To obtain more complete (although not fully complete) groundtruth motion capture, we record the human from multiple camera angles using a high-precision markerless motion capture systemfastmove_gait capable of acquiring multi-view videos and 3D skeletal points. This system comprises four synchronized ZCAM E2M4 cinema cameraszcam , each equipped with an Olympus 14–150 mm lens (fixed at 14 mm)lens , recording at 4K resolution (3840$\times$2160 pixels) with 120 FPS. To replicate recording using a typical smartphone and improving the accessibility, we additionally incorporated a smartphone view captured with a OnePlus 7phone at 1080P resolution (1920$\times$1080 pixels) and 60 FPS.
@@
Multimodal information. Existing datasets predominantly include modalities such as imagespirsiavash2014assessing , textsparmar2019and , skeletal pointscapecci2019kimore , and audioparmar2021piano , with limited exploration of other potentially valuable physiological modalities. Physiological modalities, especially EMG, which directly reflects deep muscular activity and is naturally related to movement or action quality, provide particularly relevant and deep feedback for AQA in fitness scenarios. Thus, in the FLEX dataset, we record: 1) EMG signals using a 16-channel FastMove sEMGfastmove device operating at a sampling rate of 2000 Hz. sEMG captures muscle activity. 2) Heart rate, collected with the help of a customized wearable vest, provides insight into a person’s exertion rate. 3) Respiratory rate: Exercise requires proper and procedural breathing. Our exercise procedure base and rules include breathing as an AQA criterion.
Motion capture environment and sensor placements. All devices were installed within a controlled space measuring 5m×$\times$5m×$\times$2m (as shown in Figure 4). The four cinema cameras for motion capture were fixed at each corner of the square space, while the smartphone was positioned at the main viewpoint. All four sides of the controlled space were covered with curtains to reduce environmental interference further. The placement of the sEMG sensors was adjusted based on the specific target muscles involved in each action. Each subject wore the customized vest tightly to ensure accuracy in data acquisition.
Collection process. To begin the data collection process, each subject was first informed of the actions to perform. Subsequently, a data collection staff assisted human subjects in wearing the customized vest and attaching the sEMG sensors to the target muscle regions. Notably, in contrast to data collection strategies adopted by the FLAG3Dtang2023flag3d and Egoexo-Fitnessli2024egoexo datasets, subjects in our study were required to enter the collection area individually and sequentially—to avoid observing others or receiving detailed instructions from the data collection staff. This approach prevented external factors from influencing the subjects’ performance levels. Additionally, no textual action descriptions were provided during the data collection; instead, a concise demonstration was shown briefly before collecting started, ensuring that subjects performed the actions based solely on their own experience. After the sensor attachment, the subject entered the collection area, facing the curtain between Viewpoints 1 and 2. Upon receiving the staff’s signal to begin the test, the subject independently performed the specified action 10 times, after which the staff recorded the corresponding weight data. The whole process was overseen by the authors.
Rest periods in between data collection. Considering the difficulty of performing 20 consecutive actions, particularly for the 20 novice subjects, only professional-level subjects performed all 20 consecutively. The remaining 28 subjects completed their collections separately over multiple days to ensure high-quality data, grouped by targeted muscle regions.
@@ 3.3.1 Developing Standardized Fitness Rules and Knowledge-base
Unlike competitive sports (e.g., diving, figure skating, gymnastics) with clearly defined action objectives and scoring criteria, fitness actions aim to serve diverse goals such as muscle growth, fat loss, and rehabilitation. Consequently, different individuals might adopt varying evaluation standards for the same action. To guarantee high-quality annotation for the FLEX dataset, we extensively gathered existing fitness action standards from multiple sources, including national standards, fitness associations, professional literature, fitness applications, and influential fitness content creators. We established a precise and objective annotation system through comprehensive integration and refinement.
Sources of Fitness Action Standards. Specifically, our annotation rules primarily drew upon the "National Occupational Skill Standards—Social Sports Instructor (Occupation Code: 4-13-04-01)" jointly formulated by the Ministry of Human Resources and Social Security of the People’s Republic of China and General Administration of Sport of Chinastandard2020 . Additionally, we referenced the "Occupational Competency Training Textbook for Social Sports Instructors—Fitness Coaches (with Technical Action Videos)" published by the Human Resources Development Center of the General Administration of Sport of ChinaHRDSport_Toolkit , the "Fitness and Bodybuilding Tutorial" from Beijing Sport Universitybsu2013 , and "Joe Weider’s Bodybuilding System"joe .
Developing Standardized Annotation Rules through Multi-Source Integration and Expert Review. We thoroughly reviewed all sources, meticulously documenting essential information relevant to the collected actions, emphasizing target muscles, action descriptions, and potential error types. Through rigorous comparison and integration, we distilled these insights into annotation rules that adhere to existing standards while accommodating practical training scenarios. These rules underwent rigorous verification by experts from Beijing Sport University, ensuring both theoretical rigor and enhanced consistency and practicality in annotations.
Action Segmentation and Keystep Modeling. We segmented each collected action into three distinct phases based on major limb trends: Preparation, Concentration, and Eccentric. Within each phase, the main limb actions were further subdivided into several action key-steps (AK), with integrated descriptions assigned accordingly. Combining these AKs systematically enabled the generation of textual descriptions for diverse fitness actions.
@@
Knowledge Graph Construction. Consequently, we constructed a fitness action knowledge graph (see in the Appendix) comprising actions, action keysteps (AK), error types (ET), feedback, and their interrelationships.
Scoring System for Annotation Efficiency. Furthermore, different error types were assigned distinct weights within a penalty-based scoring system, enabling annotators to efficiently identify and label errors from action clips. The final scores were obtained through cumulative penalties, ensuring both efficiency and consistency throughout the annotation process (provided in the Appendix).
Figure 5: Annotation Process. The annotators we recruited were trained according to the sources of the annotation guidelines and underwent centralized training to ensure they thoroughly understood the rules. The video data was segmented following predetermined criteria, and a two-stage annotation process was implemented to reduce annotation errors and mitigate subjective bias.
@@ 3.3.2 Annotation Process
We followed the steps for annotating our dataset.
@@
    Annotator Recruitment. We carefully recruited 16 fitness professionals and practitioners to annotate our dataset. For details on the recruitment procedure, please refer to the Appendix.
@@
    Action Segmentation. During the data segmentation phase, annotators segmented videos containing repeated actions into individual action samples according to the previous standards and guidelines and verified each sample.
@@
    Task Organization and Annotator Training. In the formal annotation phase, we organized the tasks into several work periods based on the target muscle groups of the actions and provided centralized training for 16 annotators at the beginning of each period to ensure a thorough understanding of the annotation guidelines.
@@
    Annotator Grouping by Experience. Based on their fitness experience, the annotators were divided into two groups: the first group—comprising trainers with less than three years of experience and master’s students—handled the initial annotation (Stage 1), while the second group—consisting of trainers with over three years of experience and doctoral students—was responsible for annotation verification (Stage 2).
@@
    Two-Stage Annotation & Verification Workflow. In Stage 1, each sample was initially annotated for error types by three different annotators from the first group using both video and 3D pose information. In Stage 2, these three sets of initial annotations were reviewed by three annotators from the second group. If any issues were identified, the final annotation was determined by a majority vote among these three reviewers to mitigate subjective bias.
@@
    Quality Control Measures. Authors periodically monitored the quality of the annotated data to ensure that the dataset maintained a high degree of reliability and academic value for subsequent analyses.
Dataset Statistics and Visualization. In total, we collect 7512 samples from 38 subjects doing 20 exercises, each for 10 repetitions. We have provided further dataset stats along with WordClouds and visualized samples in the Appendix.
3.4 FLEX VideoQA Dataset
We leverage LLM to generate Questions and Answers based on our Fitness Knowledge Graph and annotations related to error types, action description and feedback. These QA sets are then manually verified and corrected if needed. More details are provided in the Appendix.
@@ 4 Experiments
FLEX dataset’s Multimodal data, Fitness Action Knowledge Graph, and the fine-grained annotations not only support detailed action quality assessment but also serve as a rich resource for diverse research scenarios such as chatbot-based fitness assessment and coaching (involves video question answering) and crossmodal signal estimation. To foster future research on FLEX, we have designed and conducted multiple benchmark evaluations on the dataset to demonstrate its broad application potential.
@@ 4.1 Action Quality Assessment
Baselines. In this section, we develop the following baselines by extending the CoRe model yu2021group .
  1. 1.
    Multiview (MV): This model uses multi-view video, based on the hypothesis that combining perspectives provides a more complete view of posture and movement, enhancing AQA performance.
  2. 2.
    Multiview incorporating explicit Pose information (MV+Pose): We hypothesize that pose features can enhance the video-based AQA model. Thus this model builds on MV model and incorporates pose features extracted via an ST-GCNyan2018spatial pre-trained on NTU-RGB+Dshahroudy2016ntu .
  3. 3.
    Multiview incorporating explicit Pose information and Muscle-level Physiological information (MV+Pose+EMG): sEMG offers direct insight into muscle activity, valuable for AQA. This model extends MV+Pose by integrating sEMG. We do this by computing relative muscle contribution and concatenating it with video and pose features.
Implementation details. Provided in the Appendix. Full Codebase has been released.
Metrics. Following standard practice, we report Spearman’s rank correlation ($\rho$) and $R-l2$ metrics.
@@
AQA Model $\rho\uparrow$ $R-l2\downarrow$
Single viewyu2021group 0.8069 1.7582
@@
MV + Pose 0.9003 0.8920
MV + Pose + EMG 0.9019 0.8877
Results. For a fair comparison, we keep the architecture, loss function and training strategy same across models. Results are summarized in Figure 6. On our FLEX dataset, the single view model achieves an $\rho$ of 0.8069 and an $R\mbox{-}l2$ of 1.7582, demonstrating the efficacy of contrastive regression modeling. When multiview data is leveraged—MV model—the model attains substantial gains on both $\rho$ rises to 0.8974 and $R\mbox{-}l2$ falls to 0.9095—indicating that the second viewpoint enriches appearance information in the action space and enables more robust similarity discrimination. Further integrating skeletal point information (MV+Pose) boosts performance further ($\rho$ = 0.9003, $R\mbox{-}l2$ = 0.8920), underscoring the complementary value of high-level structural motion representations for capturing movement trends and amplitude. Finally, by introducing sEMG features (MV+Pose+EMG), the model achieves its best results, $\rho$ reaches 0.9019, and $R\mbox{-}l2$ drops to 0.8877. These findings suggest that the muscle-level control signals provided by sEMG can supplement external motion trajectories by capturing internal physiological states, thereby furnishing a more mechanistic foundation for action quality assessment.
@@ 4.2 Action Quality Understanding
We envision a future where people can use chatbots as AI-Coaches, who can understand and monitor peoples’ exercise posture and movements and provide corrective feedback on it. Vision-Language Models (VLMs) or Multimodal Large Language Models (MLLMs) are a natural choice to implement such AI-Coach because their capability to conduct multimodal communication and processing. However, out of the box, these models might not have specialized skills/knowledge to perform fitness action quality assessment. We believe our multimodal data, the fitness action knowledge graph, and the fine-grained annotations in the FLEX dataset can bridge this gap. To that end, we create FLEX-VideoQA dataset (subsection 3.4) containing fitness-based video question-answering samples. This dataset unlocks the integration of VLMs/MLLMs in AQA research.
Figure 7: Performance of Qwen2.5-VL 3B on FLEX-VideoQA dataset. BF1: BERTScore F1, SFT: Supervised Fine-Tuning. $R-l2(\times 100)$
@@
BLEU $\uparrow$ 0.0388 0.0415 0.1970
ROUGE-L $\uparrow$ 0.2012 0.2128 0.4010
METEOR $\uparrow$ 0.3049 0.3206 0.4688
CHRF++ $\uparrow$ 36.1340 39.5308 52.0005
BF1 $\uparrow$ 0.1664 0.1666 0.4390
$R-l2\downarrow$ 3.1183 3.0942 3.0592
Setup. Noting its strong performance, we choose Qwen2.5-VL-3Bbai2025qwen2 , a recently released multimodal chatbot model to develop our AI-driven Fitness Coach. Note that other VLMs can also be used. We evaluated its performance under three configurations: 1) Off-the-Shelf mode: Using the original pretrained weights. 2) Rules-based Prompting mode: Providing standardized action guidelines during the evaluation phase. 3) Supervised fine-tuning mode: Applying a LoRA strategy to fine-tune the entire model on top of the pretrained weights. Further implementation details are provided in the Appendix.
Metrics. To comprehensively assess Qwen’s outputs, we employed BLEU, ROUGE-L, METEOR, CHRF++, and BERTScore-F1, which are commonly used in LLM research, as metrics of generated text, alongside $R\mbox{-}l2$ to evaluate prediction-score accuracy.
Results. Analysis of the results in Figure 7 shows that supervised fine-tuning delivers an order-of-magnitude improvement in text-generation quality compared to the pretrained model, yet only yields marginal improvements in consistency of action scoring. The results show significant improvement in semantic alignment between fitness terminology and visual entities—not covered in mainstream video understanding datasets, but covered in our dataset. However, $R\mbox{-}l2$ decreases by less than 2%, indicating that the current fine-tuning strategy optimizes the language-generation head and contributes minimally to the regression head, which relies on fine-grained pose differences. While current VLMs still underperform specialized AQA models in precise action scoring, their ability to generate natural, professional action descriptions offers a viable path for integrating VideoQA task with AQA task.

### v3

@@ 3.2 Data Collection
Fitness actions often involve complex postures, with key joints occluded when viewed from a single angle. To obtain more complete ground-truth motion, we used a high-precision markerless motion capture system FASTMOVE (2024b) with four synchronized ZCAM E2M4 cinema cameras ZCAM (2020) (Olympus 14–150 mm lens fixed at 14 mm OLYMPUS (2022)), recording 4K video at 120 FPS. For accessibility, we also included a smartphone view from a OnePlus 7 OnePlus (2019), captured at 1080p and 60 FPS.
@@
Multimodal information. Existing datasets focus on videos Pirsiavash et al. (2014), text Parmar & Morris (2019b), skeletal points Capecci et al. (2019), or audio Parmar et al. (2021), with little attention to physiological data. EMG, in particular, captures deep muscular activity closely tied to action quality, making it highly relevant for fitness AQA and deeper feedback. Accordingly, the FLEX dataset records: 1) EMG signals of target muscle groups using a 16-channel FastMove sEMG FASTMOVE (2024a) device operating at a sampling rate of 2000 Hz. sEMG captures muscle activity. 2) Heart rate, collected with the help of a customized wearable vest, provides insight into a person’s exertion rate. 3) Respiratory rate: Exercise requires proper and procedural breathing. Our exercise procedure base and rules include breathing as an AQA criterion.
Collection process. Data collection began by informing each subject of the required actions. Staff assisted subjects in wearing the customized vest and attaching sEMG sensors to target muscles. Unlike FLAG3D Tang et al. (2023) and Egoexo-Fitness Li et al. (2024), subjects entered the collection area individually to avoid observing others or receiving detailed instructions, minimizing external influence. No textual descriptions were provided; instead, a brief demonstration preceded each session. After sensor attachment, subjects faced the curtain between Viewpoints 1 and 2 and, upon the staff’s signal, performed the action 10 times. The staff then recorded the corresponding weight data, with the entire process supervised by the authors.
@@ 3.3.1 Developing Standardized Fitness Rules and Knowledge-base
Unlike competitive sports (e.g., diving, gymnastics) with clearly defined action objectives and scoring criteria, fitness actions aim to serve diverse goals such as muscle growth, fat loss, and rehabilitation. As such, different individuals might adopt varying evaluation standards for the same action. To guarantee high-quality annotation for the FLEX dataset, we extensively gathered existing fitness action standards from multiple sources, including national standards, fitness associations, professional literature, fitness applications, and influential fitness content creators. We established a precise and objective annotation system through comprehensive integration and refinement.
Sources of Fitness Action Standards. Specifically, our annotation rules primarily drew upon the "National Occupational Skill Standards—Social Sports Instructor (Occupation Code: 4-13-04-01)" jointly formulated by the Ministry of Human Resources and Social Security of the People’s Republic of China and General Administration of Sport of China Ministry of Human Resources and Social Security of the People’s Republic of China & General Administration of Sport of China (2020). Additionally, we utilized the "Fitness and Bodybuilding Tutorial" from Beijing Sport University University (2013), "Occupational Competency Training Textbook for Social Sports Instructors—Fitness Coaches (with Technical Action Videos)" published by the Human Resources Development Center of the General Administration of Sport of China of the General Administration of Sport of China (2023), and "Joe Weider’s Bodybuilding System" Weider (1998).
Developing Standardized Annotation Rules through Multi-Source Integration and Expert Review. We thoroughly reviewed all sources, meticulously documenting essential information relevant to the collected actions, emphasizing target muscles, action descriptions, and potential error types. Through rigorous comparison and integration, we distilled these insights into annotation rules that adhere to existing standards while accommodating practical training scenarios. These rules underwent rigorous verification by experts from an advanced sports institution, ensuring both theoretical rigor and enhanced consistency and practicality in annotations.
Figure 4: Annotation Process. Annotators were trained on the provided guidelines and received centralized instruction to ensure full understanding of the rules. The video data was segmented following predetermined criteria, and a two-stage annotation process was implemented to reduce annotation errors and mitigate subjective bias.
Biomechanically-Inspired Action Phase and Keystep Modeling To guide the annotation process and ensure consistency, we developed a biomechanically-inspired framework for structuring each action. In this framework, every collected action is decomposed into three distinct phases based on dominant limb movement patterns: Preparation, Concentration, and Eccentric. Within each phase, the primary limb actions are further subdivided into finer-grained action keysteps (AKs), each paired with integrated textual descriptions. By systematically combining these AKs, we generate rich textual representations of diverse fitness actions, supporting fine-grained action analysis and structured representation learning. Building on this framework, we next introduce extensions for error identification, structured knowledge representation, and annotation efficiency.
@@
Fitness Action Knowledge Graph (FKG) Construction. We constructed a FKG (see in the Appendix B) comprising of actions, action keysteps (AK), error types (ET), feedback, and their relationships.
Scoring System for Annotation Efficiency. Furthermore, different error types were assigned distinct weights within a penalty-based scoring system, enabling annotators to efficiently identify and label errors from action clips. The final scores were obtained through cumulative penalties, ensuring both efficiency and consistency throughout the annotation process (provided in the Appendix B).
@@ 3.3.2 Annotation Process
We annotate the following information for each sample: a) Action segmentation; b) Action keysteps; c) Action Errors; d) Action Quality Assessment Scores using the following procedure.
@@
    Annotator Recruitment. We carefully recruited 16 fitness professionals and practitioners to annotate our dataset. For details on the recruitment procedure, please refer to the Appendix B.
@@
    Action Segmentation. During data segmentation, annotators split videos of repeated actions into individual samples following established guidelines and verified each sample.
@@
    Task Organization and Annotator Training. During annotation phase, tasks were organized by target muscle groups into separate work periods. At the start of each period, all 16 annotators received centralized training to ensure consistent understanding of the guidelines.
@@
    Annotator Grouping by Experience. Based on their fitness experience, annotators were organized into two groups: 1) Group 1—trainers with < 3 years of experience and master’s students—handled the initial annotation (Stage 1), while 2) Group 2—trainers with > 3 years of experience and doctoral students—was responsible for annotation verification (Stage 2).
@@
    Two-Stage Annotation & Verification Workflow. Stage 1: Each sample was first annotated for error types by three annotators from Group 1, using both video and 3D pose data. Stage 2: These annotations were then reviewed by three annotators from Group 2. If discrepancies arose, the final label was decided by majority vote among the reviewers to reduce subjective bias.
@@
    Quality Control Measures. Authors periodically monitored the quality of the annotated data to ensure that the dataset maintained a high degree of accuracy and reliability.
3.4 FLEX-VideoQA Dataset
To extend FLEX beyond action quality scoring, we introduce FLEX-VideoQA, a fine-grained video question–answering benchmark built directly on top of the dataset’s multimodal recordings and the Fitness Knowledge Graph (FKG). This benchmark is designed to test a models’ ability to perform structured, multimodal reasoning about human action—capturing not only visual appearance but also the underlying biomechanical and semantic relationships encoded in the FKG. Furthermore, in non-AQA VideoQA datasets, questions focus on scenes or coarse actions (e.g., what someone is doing, what will they do next, etc.). In contrast, we target fine-grained errors in exercises—posture, movement, and corrective feedback—which are absent from existing VideoQA datasets.
Dataset Construction. The FKG provides a rich ontology of actions, key steps, common error types, muscle groups, and corrective feedback. From this graph, we generate a large set of natural-language questions and answers by applying a mixture of rule-based templates and LLMs. Since every question is grounded in explicit graph nodes and relations, the resulting QA pairs are both semantically precise and verifiable. These QA sets are then manually verified and corrected if needed. In total, we collect 30048 QA pairs. The questions span several categories:
  • Descriptive: ask for observable attributes such as the action being performed or the current keystep (e.g., “Which exercise is shown in the video?”, “What keystep is the performer currently executing?”).
  • Relational/reasoning: require linking two or more graph entities (e.g., “Which muscle group is primarily activated during the descent phase of a squat?” or “Which error in the setup step leads to a lower-back penalty?”).
  • Temporal: probe understanding of sequential dependencies (e.g., “Which key step follows the liftoff phase?”).
  • Causal/feedback: ask for corrective advice given an observed error (e.g., “What feedback should be given if the knees cave inward during the ascent?”).
Comprehensive multimodal reasoning platform. By transforming the FKG’s structured annotations into a large set of graph-grounded question–answer pairs, FLEX-VideoQA evolves FLEX from a dataset for action quality assessment into a comprehensive multimodal reasoning platform. It provides a rigorous testbed for training and evaluating VLMs that must not only recognize human actions but also reason over the hierarchical relationships among actions, key steps, muscle activations, common errors, and corrective feedback. The supervised fine-tuning experiments underscore the value of structured graph supervision for enabling VLMs to acquire representations of action quality and underlying physiology that go well beyond surface-level video understanding.
@@ 4 Experiments
FLEX dataset’s Multimodal data, Fitness Knowledge Graph, and the fine-grained annotations not only support detailed action quality assessment but also serve as a rich resource for diverse research scenarios such as chatbot-based fitness assessment and coaching (involves VideoQA) and crossmodal signal estimation. To foster future research on FLEX, we have designed and conducted multiple benchmark evaluations on the dataset to demonstrate its broad application potential.
@@ 4.1 Action Quality Assessment
Baselines. We first investigate the impact of various modalities individually and their combination for the task of AQA. Given the wide adoption and strong performance of: a) CoRe and TPT on video-based AQA, we consider them for video-based finegrained action modeling; b) STGCN and SkateFormer on pose-based action recognition, we consider them for pose-based finegrained action modeling. For EMG modality, we develop a comparative model—computing relative muscle contribution and concatenating it with video and pose features. Based on these models, we develop the following baselines. 1) Unimodal models: We compared model performance across modalities, thereby isolating each modality’s independent contribution. 2) Multiview: This model uses multi-view video, based on the hypothesis that combining perspectives provides a more complete view of posture and movement, enhancing AQA performance. 3) Multiview incorporating explicit Pose information: We hypothesize that pose features can enhance the video-based AQA model. Thus this model builds on MV model and incorporates pose features extracted via an ST-GCN Yan et al. (2018) or SkateFormer Do & Kim (2024). 4) Multiview incorporating explicit Pose information and Muscle-level Physiological information: sEMG offers direct insight into muscle activity, valuable for AQA. This model extends MV+Pose by integrating sEMG. We do this by computing relative muscle contribution and concatenating it with video and pose features. Implementation details. Provided in the Appendix C. Metrics. Following standard practice, we report Spearman’s rank correlation ($\rho$) and Rl2R-l2 metrics.
@@
AQA Model $\rho\uparrow$ $R-\ell_{2}\downarrow$
CoRe Yu et al. (2021) 0.8069 1.7582
TPT Bai et al. (2022) 0.8111 1.5464
ST-GCN Yan et al. (2018) 0.8528 1.3273
SkateFormer Do & Kim (2024) 0.8145 1.6645
EMG 0.3191 5.1019
@@
MV + SkateFormer 0.8982 0.9064
MV + ST-GCN 0.9003 0.8920
MV + SkateFormer + EMG 0.8996 0.9017
MV + ST-GCN + EMG 0.9019 0.8877
Results. Results of AQA models are summarized in Figure 5. In unimodal settings, TPT and CoRe perform comparably; we adopt CoRe for subsequent experiments owing to its lower computational demands and more streamlined design. Pose alone—implemented via ST-GCN & SkateFormer—outperforms single view video, as skeletal points encode fine-grained kinematics. sEMG alone yields lower performance, yet provides complementary information by capturing muscle activation and motor effort not visible in video or pose. This complementarity is confirmed in multimodal configurations: adding pose to multiview video improves performance, indicating that structural pose features enrich visual dynamics, while further integrating sEMG delivers the best overall results ($\rho$ = 0.9019, $R-\ell_{2}$ = 0.8877). By explicitly modeling external kinematics (pose) and internal physiological states (sEMG), the multimodal framework furnishes a more comprehensive and mechanistic basis for action quality assessment. For more detailed analysis, please refer to the Appendix C.
@@ 4.2 Action Quality Understanding
We envision a future where people can use chatbots as AI-Coaches, who can understand and monitor people’s exercise posture and actions and provide corrective feedback on it. Vision-Language Models (VLMs) or Multimodal Large Language Models (MLLMs) are a natural choice to implement such AI-Coach because of their capability to conduct multimodal communication and processing. However, out of the box, these models might not have specialized skills/knowledge to perform fitness action quality assessment. We believe our multimodal data, the fitness action knowledge graph, and the fine-grained annotations in the FLEX dataset can bridge this gap. To that end, we create FLEX-VideoQA dataset (Sec. 3.4) containing fitness-based video question-answering samples. This dataset unlocks the integration of VLMs/MLLMs in AQA research.
Setup. To comprehensively evaluate the performance of existing VLMs on the FLEX-VideoQA dataset, we selected several recently released SOTA models, including MiniCPM-O-2.6-8B Yao et al. (2024), InternVL3-2B Zhu et al. (2025), and Qwen2.5-VL-3B Bai et al. (2025). We evaluated their performance under three configurations: 1) Off-the-Shelf mode: Using the original pretrained weights. 2) Rules-based Prompting mode: Providing standardized action guidelines during the evaluation phase. 3) Supervised fine-tuning mode: Applying a LoRA strategy to fine-tune the entire model on top of the pretrained weights. Further implementation details are provided in Appendix C.
Metrics. To comprehensively assess VLMs’ outputs, we employed a wide range of commonly used metrics in VideoQA research (BLEU, ROUGE-L, METEOR, CHRF++, and BERTScore‑F1), as metrics of generated text, alongside $R-\ell_{2}$ to evaluate score-prediction accuracy.
Figure 6: Performance of VLMs on FLEX-VideoQA dataset. BF1: BERTScore F1, SFT: Supervised Fine-Tuning. $R-\ell_{2}(\times 100)$
@@
InternVL MiniCPM Qwen InternVL MiniCPM Qwen Qwen
BLEU $\uparrow$ 0.0003 0.0154 0.0388 0.0007 0.0312 0.0415 0.1970
ROUGE-L $\uparrow$ 0.1689 0.2902 0.2012 0.2002 0.3102 0.2128 0.4010
METEOR $\uparrow$ 0.1405 0.3024 0.3049 0.1736 0.3345 0.3206 0.4688
CHRF++ $\uparrow$ 8.7228 20.4180 36.1340 17.3078 28.4865 39.5308 52.0005
BF1 $\uparrow$ 0.1017 0.2535 0.1664 0.1550 0.2743 0.1666 0.4390
$R-\ell_{2}\downarrow$ 3.9900 10.5539 3.1183 3.2757 8.6448 3.0942 3.0592
Results. Analysis of the results in Figure 6 shows that supervised fine-tuning delivers an order-of-magnitude improvement in text-generation quality compared to the pretrained model, yet only yields marginal improvements in consistency of action scoring. The results show significant improvement in semantic alignment between fitness terminology and visual entities—not covered in mainstream video understanding datasets, but covered in our dataset. Importantly, across all three evaluation configurations, Qwen consistently achieves the best overall performance, even surpassing larger models such as MiniCPM. In addition, prompt engineering brings moderate but noticeable improvements over the pretrained baseline, further validating the benefit of simple adaptation strategies before full fine-tuning. However, $R-\ell_{2}$ decreases by less than 2%, indicating that the current fine-tuning strategy optimizes the language‐generation head and contributes minimally to the regression head, which relies on fine-grained pose differences. While current VLMs still underperform specialized AQA models in precise action scoring, their ability to generate natural, professional action descriptions offers a viable path for integrating VideoQA task with AQA task. For more detailed analysis, please refer to the Appendix C.